# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 7.4.1 The phase diagram of the three-dimensional ${\mathbb{Z}}_{2}$-gauge Higgs model
A sketch of the phase diagram is shown in Fig. 15. For $K\to\infty$ an Ising transition occurs at [316] $J_{\rm Is}=0.221654626(5)$. By duality, in the pure ${\mathbb{Z}}_{2}$ gauge model a transition occurs in the corresponding point, $J=0$ and $K_{c}=-{1\over 2}{\rm ln}\,{\rm tanh}\,J_{\rm Is}=0.761413292(11)$. Two Ising-like continuous transition lines, related by the duality transformation (78), start from these points [41, 342] (the argument is analogous to that discussed in Sec. 7.2.3) and intersect on the self-dual line [343, 174] at $K_{\star}=0.7525(1)$, $J_{\star}=0.22578(5)$. Note that the transitions along both lines are Ising-like transitions and that in both cases there is no global symmetry. Nonetheless, as we shall discuss, the two lines have different nature: the line that starts at $J=0$ is a topological transition line, while the line that starts at $K=\infty$ is an Ising× line, i.e., it admits an Ising order parameter that can only be identified once a proper gauge fixing is introduced.
In the absence of a gauge fixing, only the cumulants of the energy density (see Sec. A.3.4) can be used to characterize these transitions in a gauge-invariant way  [343, 174] (note that the tensor $R^{ab}$ defined in Eq. (70) trivially vanishes for $N=1$). Finally, some numerical studies [344, 345, 346] have provided evidence of first-order transitions along the self-dual line, in the relatively small interval starting from the MCP and ending at $J_{\star}\approx 0.258$ and $\kappa_{\star}\approx 0.688$. This endpoint is expected to correspond to a continuous transition belonging to the Ising universality class (see Ref. [343] for numerical checks).
@@
Since the first-order transition line is limited to a finite interval along the self-dual line, only two thermodynamic phases exist, see Fig. 15. For small $J$ and large $K$ there is a topological deconfined phase. The remaining part of the phase diagram corresponds to a single phase that extends from the disordered small-$J,K$ region to the whole large-$J$ region. In particular, no phase transition occurs along the line $K=0$, where the model becomes trivial, as can be seen in the unitary gauge $s_{\bm{x}}=1$. We may only distinguish two different regimes: a Higgs-like regime in the large $J$ and $K$ region, and a confined regime in the small $J$ and $K$ region (see, e.g., [350]). However, these two regions are connected by continuous paths in the phase diagram, along which all local gauge-invariant operators (more precisely the gauge invariant operators depending just on a finite number of $s_{\bm{x}}$ and $\sigma_{{\bm{x}},\mu}$ variables) are analytic functions of the Hamiltonian parameters [41, 342].
@@ 7.4.2 Order parameters driving the Ising× transitions
Along the two transition lines that start at $J=0$ and $K=\infty$ and are related by duality, thermal observables behave as in the Ising universality class. However, as already discussed in Sec. 3.5.2, the corresponding universality classes should be considered as different. Even if there is no ${\mathbb{Z}}_{2}$ global symmetry, the transitions along the line that starts at $K=\infty$ are Ising× transitions and admit, as we shall describe below, an effective description in terms of a standard LGW theory with ${\mathbb{Z}}_{2}$ global invariance, once an appropriate gauge fixing has been introduced. On the other hand, along the line that starts at $J=0$, transitions are topological as in the pure ${\mathbb{Z}}_{2}$ gauge model.  [^37] Note that the topological transitions along the transition line starting from the $J=0$ critical point cannot be characterized by a change of size behavior of the Wilson loops, as that of the 3D ${\mathbb{Z}}_{2}$ gauge model, see Sec. 3.5.2, because the Wilson loop satisfies the perimeter law for any $J>0$, due to the screening of the site spin variables. However, a more general interpretation of these topological transitions is discussed in Ref. [20] within quantum formulations, see also Refs. [351, 352, 353, 354, 355, 356, 357], as a topological transition between a small-$K$ trivial phase and a large-$K$ phase with topological order, which would respectively correspond to the confined and deconfined phases at $J=0$. We also mention that these transitions are sometimes referred to as Ising transitions; see, e.g., Refs. [345, 317, 343, 174]. As discussed in Sec. 3.5.2 we will call them Ising gauge transitions.
@@
To verify the Ising× critical behavior, Ref. [232] studied the ${\mathbb{Z}}_{2}$-gauge Higgs model supplemented with the stochastic gauge fixing introduced in Sec. 7.3, using the ancillary Hamiltonian $H_{w}$ defined in Eq. (76). Numerical FSS analyses [232] along the large-$K$ transition line show that the critical behavior of the correlation function $G_{V}({\bm{x}},{\bm{y}})$ defined in Eq. (75) is the same as in the standard Ising model, with the large-$J$, large-$K$ region corresponding to the magnetized phase. Therefore, as in multicomponent models the stochastic gauge fixing allows us to identify the universal gauge-dependent spin correlations which characterize standard Ising transitions. The same conclusions were reached in Ref. [358], in what was called Landau gauge (see footnote 33), which essentially corresponds to taking the limit $\gamma\to\infty$ in the stochastic gauge fixing approach (but note that the multiplicity of the minima, the so-called Gribov copies, could not be taken into account). The same analysis, using the stochastic gauge fixing, was also performed along the small-$J$ transition line. In this case the correlation function $G_{V}({\bm{x}},{\bm{y}})$ is not critical at the transition. As expected, these transitions are not LGW× transitions.
@@
In the stochastic gauge-fixed theory, the spin $s_{\bm{x}}$ magnetizes along the Ising× transition line, i.e., there is a ferromagnetic phase in the large-$J$, large-$K$ Higgs phase. On the other hand, for small values of $J$, $s_{\bm{x}}$ is always disordered. This implies the presence of a transition line in the gauge-fixed theory that separates the single phase [41], in which the confining and the Higgs regime both occur. The line most probably starts at the CEP, see Fig. 15, and necessarily ends on the $K=0$ axis.383838 For $K=0$ the gauge-fixed model is equivalent to the so-called ±$\pm J$ Ising model [151, 326, 327, 328, 329, 330, 331, 332, 333, 334, 335, 336, 337, 338, 339, 340], with spatially-independent random links (it is immediate to verify it in the so called unitary gauge $s_{\bm{x}}=1$, see footnote 34). In the notation of Refs. [334, 335, 336], the probability $p$ of the random link variable $J_{\rm xy}$ ($\sigma_{{\bm{x}},\mu}$ in our model in the unitary gauge) is $P(J_{\rm xy})=p\delta(J_{\rm xy}-1)+(1-p)\delta(J_{\rm xy}+1)$ with $p=e^{J}/(e^{J}+e^{-J})$. The temperature $T$ in the ±$\pm J$ Ising model corresponds to $T=\gamma^{-1}$, so the line $J=\gamma$ corresponds to the so-called Nishimori line [331]. Neglecting the small reentrance of the ferromagnetic-glassy transition line—see Refs. [334, 335, 336] for a discussion of the phase diagram of the ±$\pm J$ Ising model—,the ancillary model has a large-$\gamma$ ferromagnetic phase only for $J>1/T_{\rm N}=J_{N}=0.5991(1)$ ($T_{N}$ is the Nishimori temperature). Moreover, in the ferromagnetic phase the correlation function $G_{V}$ orders (note that the ancillary model has a ferromagnetic phase also close to the small-$J$ topological line, but here $G_{V}$ does not order). This implies that the transition line ends in $K=0$, $J=J_{N}$ [359]. This is confirmed by numerical analyses [359], which observe continuous transitions in the small-$K$ region, on the left of the first-order transition line, see Fig. 15. This line, which separates the Higgs phase, in which the spin variables $s_{\bm{x}}$ magnetize, from the confined phase, is related to the specific form of the ancillary model and therefore does not have a direct physical significance. However, it signals that the thermodynamic description does not provide the full picture.
@@
To distinguish the confinement/deconfinement properties of the model, several nonlocal order parameters have been proposed, see, e.g., Refs.[351, 352, 353, 360, 354, 355, 356, 357, 20, 348]. Note that the size dependence of Wilson loops does not provide a good order parameter. In the presence of dynamical matter fields transforming in the fundamental representation, charge screening is present and Wilson loops always scale with the perimeter law. Only for $J=0$ does the transition at $K=K_{\mathbb{Z}_{2}}$ separate regions characterized by the area law and the perimeter law. A possible order parameter is the so-called Fredenhagen-Marcu [45, 361] order parameter, which is defined by using a gauge-invariant two-point function built by joining two spins $s_{\bm{x}}$ with a staple-shaped string of gauge fields, normalized by the square root of a Wilson loop. The behavior of this order parameter in the $\mathbb{Z}_{2}$-gauge Higgs model has been studied in Refs. [348, 362]. It correctly identifies the confinement/deconfinement transition in the model and its Ising× universality class. A different order parameter has been advocated in Ref. [363] (see also Ref. [360]), defined by using a properly normalized correlator between a Wilson and ’t Hooft loop, and studied numerically in Ref. [362]. Another order parameter, whose definition is strictly related with the strong-coupling surface formulation of the $\mathbb{Z}_{2}$-gauge Higgs model (see, e.g., [364, 343]), has been proposed and numerically tested in Ref. [365]. The possibility that the crossover between the confinement and the Higgs regime could induce superficial phase transitions in quantum many-body systems has instead been investigated in Refs. [366, 367].
7.4.3 Multicritical behavior at the meeting point of the transition lines
@@
The first-order and the two continuous Ising-like transition lines intersect in one point located along the self-dual line, giving rise to a multicritical point (MCP) [174, 343, 345]. (see A.4 for a brief review of the main universal features of the scaling behavior close to a MCP). Critical exponents have been numerically computed at the MCP [174, 343, 347, 348], obtaining results consistent with a multicritical XY behavior.
@@
The multicritical model defined in Sec. 2.2.4 has been extensively studied [170, 171, 95, 173, 250, 174]. In particular, for $N_{1}=N_{2}=1$ it admits a bicritical point (see the left panel of Fig. 1) with a critical behavior controlled by a stable fixed point with enlarged XY global symmetry. The main conjecture put forward in Ref. [174] is that this stable XY FP is the relevant one that controls the multicritical XY behavior in the $\mathbb{Z}_{2}$-gauge Higgs theory. In this scenario, the two relevant operators at the MCP are the quadratic spin-2 and spin-0 field combinations, whose RG dimensions are [173, 250] $y_{1}=1.7639(11)$ and [289] $y_{2}=1/\nu_{\rm XY}=1.4888(2)$ respectively. The numerical estimates [343, 174, 347, 348] of $y_{1}$ and $y_{2}$ are in good agreement (within errors of the order of 1%) with the XY more accurate estimates reported above. To give an idea of the accuracy of the multicritical XY scenario, in Fig. 16 we show the FSS behavior of some particular third- and fourth-order energy cumulants using the XY critical exponent $y_{2}=1/\nu_{\rm XY}$. We should note that the multicritical XY scenario is based on several plausible, but unproved assumptions, the main one being the existence of a (dual) local order parameter for the topological transitions. We should also mention that this assumption is still controversial: Refs. [343, 365, 347] consider it implausible, and interpret the numerical agreement between Monte Carlo results and XY estimates as a mere coincidence.
@@ 7.5 Other models with discrete symmetries
In Sections 5 and 6 we discussed the phase diagram and critical properties of U(1) gauge theories with global U($N$) symmetry, while in this section we considered gauge models with discrete $\mathbb{Z}_{2}$ gauge symmetry and O($N$) global symmetry. We now present some additional models with discrete gauge and/or global symmetry, in which the gauge and/or scalar fields take only a discrete set of values. These models can be considered as approximations of models with continuous groups. However, classical theorems by Jordan and Turing [368] show that the only continuous groups of transformations that can be approximated with arbitrary precision by using their discrete subgroups are the Abelian ones. Thus, in the compact case, the only possibility is replacing the U(1) group with the cyclic finite group ${\mathbb{Z}}_{N}$. In the noncompact case, the additive group ${\mathbb{R}}$ can be replaced by the discrete, but still infinite, additive group ${\mathbb{Z}}$. As we shall see, in some cases these discretized models undergo transitions that are analogous to those occurring in models with continuous fields and symmetry groups, with an enlargement of the global and/or local gauge symmetry.
@@ 7.5.1 Compact models with ${\mathbb{Z}}_{\cal N}$ local symmetry
We consider a model with global $\mathbb{Z}_{q}$ symmetry and local $\mathbb{Z}_{\mathcal{N}}$ gauge symmetry. We assume that $q$ is an integer multiple of $\mathcal{N}$ and we define $p=q/\mathcal{N}$. The scalar fields are complex phases $z_{\bm{x}}$, which take $q$ different values ($z_{\bm{x}}=\exp(2\pi im/q)$, $m=1,\ldots q-1$), while gauge fields are complex phases $\lambda_{{\bm{x}},\nu}=\exp(2\pi ni/\mathcal{N})$, $n=0,\ldots,{\mathcal{N}}-1$. The Hamiltonian and the partition function are given by
@@
For $J\to\infty$ no transition is expected (fixing the global and local invariance we can set $z_{\bm{x}}=\lambda_{{\bm{x}},\mu}=1$ on all links), while for $\kappa=0$ one can show, using universality arguments, the transition to be a LGW transition in the $\mathbb{Z}_{p}$ universality class, see Ref. [249] for more details. Thus, no transition occurs for $p=1$. The model undergoes an Ising and XY transitions for $p=2$ and $p\geq 5$, respectively, and a first-order transition for $p=3$. An XY transition is also expected for $p=4$. Indeed, in the latter case the model obtained for $\kappa=0$ is a generic model with ${\mathbb{Z}}_{4}$ global symmetry, and thus it should have an XY transition, as discussed in Sec. 3.5.2 and, in particular, in footnote 16. The order parameter at these transitions is the gauge-invariant operator $s_{\bm{x}}^{\mathcal{N}}$. This quantity transforms nontrivially only under the ${\mathbb{Z}}_{p}$ subgroup of the global invariance group, so the effective description is given by the LGW Lagrangian (21) with $Q=p$.
From the discussion of the limiting cases it follows that the phase diagram for $p>1$ has the form sketched in Fig. 17, while the phase diagram for $p=1$ is qualitatively similar to the one reported in Fig. 15, with no transition for $\kappa=0$. It is reasonable to assume the transitions along the DD-O line to be analogous to those occurring for $\kappa=0$, while the transitions along the DD-DO are expected to be topological transitions as for $J=0$. Again, the case ${\cal N}=4$ should be discussed separately. Indeed, while the transition for $J=0$ is an Ising gauge transition, for $J\not=0$ we expect an XY gauge phase transition (see the discussion in Sec. 3.5.2 and, in particular, footnote 16).
@@
It is interesting to note that the model with $\mathcal{N}=2$ and any $p\geq 5$ has the same phase diagram and critical behaviors as the $\mathbb{Z}_{2}$-gauge $N$-vector model with $N=2$, which can be obtained by performing a naive limit $q\to\infty$ of the Hamiltonian (81).404040 For $\mathcal{N}=2$ the correlation function $G_{\mathcal{N}}({\bm{x}},{\bm{y}})$ defined in Eq. (82) is equal to $2G_{R}({\bm{x}},{\bm{y}})$ [see Eq. (71)] for $N=2$, if we identify ${\bm{s}}_{\bm{x}}=(\mathrm{Re}\,z_{\bm{x}},\mathrm{Im}\,z_{\bm{x}})$ in Eq. (70). Therefore, for $q\to\infty$ the function $G_{\mathcal{N}}({\bm{x}},{\bm{y}})$ is equivalent to $G_{R}({\bm{x}},{\bm{y}})$ in the $\mathbb{Z}_{2}$-gauge XY model. Thus, for $p\geq 5$ we observe a symmetry enlargement of the global symmetry along the DD-O and DO-O lines, while, for ${\cal N}\geq 4$, correlation functions display an enlarged gauge symmetry along the topological DD-DO transition line.

### v3

@@ 7.4.1 The phase diagram of the three-dimensional ${\mathbb{Z}}_{2}$-gauge Higgs model
A sketch of the phase diagram is shown in Fig. 15. For $K\to\infty$ an Ising transition occurs at [162] $J_{\rm Is}=0.221654626(5)$. By duality, in the pure ${\mathbb{Z}}_{2}$ gauge model a transition occurs in the corresponding point, $J=0$ and $K_{c}=-{1\over 2}{\rm ln}\,{\rm tanh}\,J_{\rm Is}=0.761413292(11)$. Two Ising continuous transition lines, related by the duality transformation (81), start from these points [41, 375] (the argument is analogous to that discussed in Sec. 7.2.3) and intersect on the self-dual line [376, 205] at $K_{\star}=0.7525(1)$, $J_{\star}=0.22578(5)$. Note that the transitions along both lines are Ising-like transitions and that in both cases there is no global symmetry. Nonetheless, as we shall discuss, the two lines have different nature: the line that starts at $J=0$ is a topological transition line, while the line that starts at $K=\infty$ is an Ising× line, i.e., it admits an Ising order parameter that can only be identified once a proper gauge fixing is introduced. Note that, in the absence of a gauge fixing, only the cumulants of the energy density (see Sec. A.3.2 for definitions) can be used to characterize these transitions in a gauge-invariant way  [376, 205] (note that the tensor $R^{ab}$ defined in Eq. (73) trivially vanishes for $N=1$).
Finally, numerical studies [377, 378, 379] have provided evidence of first-order transitions along the self-dual line, in the relatively small interval starting from the MCP and ending at $J_{\star}\approx 0.258$ and $\kappa_{\star}\approx 0.688$. This endpoint is expected to correspond to a continuous transition belonging to the Ising universality class (see Ref. [376] for a numerical study).
@@
Since the first-order transition line is limited to a finite interval along the self-dual line, only two thermodynamic phases exist, see Fig. 15. For small $J$ and large $K$ there is a topological deconfined phase. The remaining part of the phase diagram corresponds to a single phase that extends from the disordered small-$J,K$ region to the whole large-$J$ region. In particular, no phase transition occurs along the line $K=0$, where the model becomes trivial, as can be seen in the unitary gauge $s_{\bm{x}}=1$. We may only distinguish two different regimes: a Higgs-like regime in the large $J$ and $K$ region, and a confined regime in the small $J$ and $K$ region (see, e.g., Ref. [383]). However, these two regions are connected by continuous paths in the phase diagram, along which all local gauge-invariant operators (more precisely the gauge invariant operators depending on a finite number of $s_{\bm{x}}$ and $\sigma_{{\bm{x}},\mu}$ variables) are analytic functions of the Hamiltonian parameters [41, 375].
@@ 7.4.2 Order parameters driving the Ising× transitions
Along the two transition lines that start at $J=0$ and $K=\infty$ and are related by duality, thermal observables behave as in the Ising universality class. However, as already discussed in Sec. 3.5.2, the corresponding universality classes are different. Even if there is no ${\mathbb{Z}}_{2}$ global symmetry, the transitions along the line that starts at $K=\infty$ are Ising× transitions and admit, as we shall describe below, an effective description in terms of a standard LGW theory with ${\mathbb{Z}}_{2}$ global invariance, once an appropriate gauge fixing has been introduced. On the other hand, along the line that starts at $J=0$, transitions are topological as in the pure ${\mathbb{Z}}_{2}$ gauge model. [^38] Note that the topological transitions along the transition line starting from the $J=0$ critical point cannot be characterized by the size behavior of the Wilson loops, at variance with what occurs in the 3D ${\mathbb{Z}}_{2}$ gauge model, see Sec. 3.5.2, because Wilson loops satisfy the perimeter law for any $J>0$, due to the screening of the site spin variables. A more general interpretation of these topological transitions is discussed in Ref. [20] in a quantum setting, see also Refs. [384, 385, 386, 387, 388, 389, 390]. They correspond to topological transitions between a small-$K$ trivial phase and a large-$K$ phase with topological order. These two phases correspond to the confined and deconfined phases at $J=0$, respectively. We also mention that these transitions are sometimes referred to as Ising transitions; see, e.g., Refs. [378, 349, 376, 205]. As discussed in Sec. 3.5.2, we name them Ising gauge transitions.
@@
To verify the Ising× critical behavior, Ref. [263] studied the ${\mathbb{Z}}_{2}$-gauge Higgs model with the stochastic gauge fixing introduced in Sec. 7.3, using the ancillary Hamiltonian $H_{w}$ defined in Eq. (79). Numerical FSS analyses [263] along the large-$K$ transition line show that the critical behavior of the correlation function $G_{V}({\bm{x}},{\bm{y}})$ defined in Eq. (78) is the same as in the standard Ising model, with the large-$J$, large-$K$ region corresponding to the magnetized phase. Therefore, as in multicomponent models the stochastic gauge fixing allows us to identify the universal gauge-dependent spin correlations which characterize standard Ising transitions. The same conclusions were reached in Ref. [391], in what was called Landau gauge (see footnote 34), which essentially corresponds to taking the limit $\gamma\to\infty$ in the stochastic gauge fixing approach (but note that the multiplicity of the minima, the so-called Gribov copies, could not be taken into account). The same analysis, using the stochastic gauge fixing, was also performed along the small-$J$ transition line. In this case the correlation function $G_{V}({\bm{x}},{\bm{y}})$ is not critical at the transition. As expected, these transitions are not LGW× transitions.
@@
In the stochastic gauge-fixed theory, the spin $s_{\bm{x}}$ magnetizes along the Ising× transition line, i.e., there is a ferromagnetic phase in the large-$J$, large-$K$ Higgs phase. On the other hand, for small values of $J$, $s_{\bm{x}}$ is always disordered. This implies the presence of an additional transition line in the gauge-fixed theory,which separates the single phase [41] in which the confining and the Higgs regime both occur. The line most probably starts at the CEP, see Fig. 15, and necessarily ends on the $K=0$ axis. [^39] For $K=0$ the gauge-fixed model is equivalent to the so-called ±$\pm J$ Ising model [180, 358, 359, 360, 361, 362, 363, 364, 365, 366, 367, 368, 369, 370, 371, 372], with spatially-independent random links (it is immediate to verify it in the so-called unitary gauge $s_{\bm{x}}=1$, see footnote 35). In the notation of Refs. [366, 367, 368], the probability $p$ of the random link variable $J_{\rm xy}$ ($\sigma_{{\bm{x}},\mu}$ in our model in the unitary gauge) is $P(J_{\rm xy})=p\delta(J_{\rm xy}-1)+(1-p)\delta(J_{\rm xy}+1)$ with $p=e^{J}/(e^{J}+e^{-J})$. The temperature $T$ in the ±$\pm J$ Ising model corresponds to $T=\gamma^{-1}$, so the line $J=\gamma$ corresponds to the so-called Nishimori line [363]. Neglecting the small reentrance of the ferromagnetic-glassy transition line—see Refs. [366, 367, 368] for a discussion of the phase diagram of the ±$\pm J$ Ising model—,the ancillary model has a large-$\gamma$ ferromagnetic phase only for $J>1/T_{\rm N}=J_{N}=0.5991(1)$ ($T_{N}$ is the Nishimori temperature). Moreover, in the ferromagnetic phase the correlation function $G_{V}$ orders (note that the ancillary model has a ferromagnetic phase also close to the small-$J$ topological line, but here $G_{V}$ does not order). This implies that the transition line ends in $K=0$, $J=J_{N}$. This is confirmed by numerical analyses, which observe continuous transitions in the small-$K$ region, on the left of the first-order transition line, see Fig. 15. This line, which separates the Higgs regime, in which the gauge-fixed spin correlations (in the generalized unitary gauge they are equivalent to correlations of the ancillary fields) magnetize, from the confined regime, is related to the specific form of the ancillary model and therefore does not have a direct physical significance. However, it signals that the thermodynamic description does not provide the full picture.
@@
To distinguish the confinement/deconfinement properties of the model, several nonlocal order parameters have been proposed, see, e.g., Refs.[384, 385, 386, 392, 387, 388, 389, 390, 20, 381]. Note that the size dependence of Wilson loops does not provide a good order parameter. In the presence of dynamical matter fields transforming in the fundamental representation, charge screening is present and Wilson loops always scale with the perimeter law. Only for $J=0$ does the transition at $K=K_{\mathbb{Z}_{2}}$ separate regions characterized by the area law and the perimeter law. A possible order parameter is the so-called Fredenhagen-Marcu [45, 393] order parameter, which is defined by using a gauge-invariant two-point function built by joining two spins $s_{\bm{x}}$ with a staple-shaped string of gauge fields, normalized by the square root of a Wilson loop. The behavior of this order parameter in the $\mathbb{Z}_{2}$-gauge Higgs model has been studied in Refs. [381, 394]. It correctly identifies the confinement/deconfinement transition in the model and its Ising× universality class. A different order parameter has been advocated in Ref. [395] (see also Ref. [392]), defined by using a properly normalized correlator between a Wilson and ’t Hooft loop, and studied numerically in Ref. [394]. Another order parameter, whose definition is strictly related with the strong-coupling surface formulation of the $\mathbb{Z}_{2}$-gauge Higgs model (see, e.g., Refs. [396, 376]), has been proposed and numerically tested in Ref. [397]. The possibility that the crossover between the confinement and the Higgs regime could induce superficial phase transitions in quantum many-body systems has instead been investigated in Refs. [398, 399].
7.4.3 Multicritical behavior at the intersection point of the transition lines
@@
The first-order and the two continuous Ising transition lines intersect in a multicritical point (MCP) located along the self-dual line [205, 376, 378]. [^40] See A.4 for a brief review of the main universal features of the scaling behavior close to a MCP. Critical exponents have been numerically computed at the MCP [205, 376, 380, 381], obtaining results consistent with a multicritical XY behavior.
@@
The multicritical model defined in Sec. 2.2.4 has been extensively studied [201, 202, 96, 204, 157, 205]. In particular, for $N_{1}=N_{2}=1$ it admits a bicritical point (see the left panel of Fig. 1) with a critical behavior controlled by a stable fixed point with enlarged O(2) global symmetry. The main conjecture put forward in Ref. [205] is that this stable multicritical XY FP is the one that controls the multicritical XY behavior in the $\mathbb{Z}_{2}$-gauge Higgs theory. In this scenario, the two relevant operators at the MCP are the quadratic spin-2 and spin-0 field combinations, whose RG dimensions are [204, 157] $y_{1}=1.7639(11)$ and [154] $y_{2}=1/\nu_{\rm XY}=1.4888(2)$ respectively. The numerical estimates [376, 205, 380, 381] of $y_{1}$ and $y_{2}$ are in good agreement (errors are of the order of 1%) with the more accurate XY estimates reported above. To give an idea of the accuracy of the multicritical XY scenario, in Fig. 16 we show the FSS behavior of some particular third- and fourth-order energy cumulants using the XY critical exponent $y_{2}=1/\nu_{\rm XY}$. The agreement with the multicritical XY scenario is globally very good and strongly supports it. However, we also note that the multicritical XY theory is based on some unproved assumptions, the main one being the existence of a (dual) local order parameter for the topological transitions. We should also mention that this assumption is still controversial: Refs. [376, 397, 380] consider it implausible, and interpret the agreement between numerical results and XY estimates as a mere coincidence.
@@ 7.5 Other models with discrete symmetries
In Secs. 5 and 6 we discussed the phase diagram and critical properties of U(1) gauge theories with global U($N$) symmetry, while in this section we considered gauge models with discrete $\mathbb{Z}_{2}$ gauge symmetry and O($N$) global symmetry. We now present some additional models with discrete gauge and/or global symmetry, in which the gauge and/or the scalar fields take only a discrete set of values. These models can be considered as approximations of models with continuous groups. However, classical theorems by Jordan and Turing [400] show that the only continuous groups of transformations that can be approximated with arbitrary precision by using their discrete subgroups are the Abelian ones. Thus, in the compact case, the only possibility is replacing the U(1) group with the cyclic finite group ${\mathbb{Z}}_{N}$. In the noncompact case, the additive group ${\mathbb{R}}$ can be replaced by the discrete, but still infinite, additive group ${\mathbb{Z}}$. As we shall see, in some cases these discretized models undergo transitions that are analogous to those occurring in models with continuous fields and symmetry groups, with an enlargement of the global and/or local gauge symmetry.
@@ 7.5.1 Compact models with ${\mathbb{Z}}_{\cal N}$ local symmetry
We consider a model with global $\mathbb{Z}_{q}$ symmetry and local $\mathbb{Z}_{\mathcal{N}}$ gauge symmetry. We assume that $q$ is an integer multiple of $\mathcal{N}$ and we define $p=q/\mathcal{N}$. The scalar fields are complex phases $z_{\bm{x}}$, which take $q$ different values, $z_{\bm{x}}=\exp(2\pi im/q)$ with $m=1,\ldots q-1$, while gauge fields are complex phases $\lambda_{{\bm{x}},\nu}=\exp(2\pi ni/\mathcal{N})$, $n=0,\ldots,{\mathcal{N}}-1$. The Hamiltonian and the partition function are given by
@@
For $J\to\infty$ no transition is expected (fixing the global and local invariance we can set $z_{\bm{x}}=\lambda_{{\bm{x}},\mu}=1$ on all links), while for $\kappa=0$ one can show, using universality arguments, the transition to be a LGW transition in the $\mathbb{Z}_{p}$ universality class, see Ref. [282] for more details. Thus, no transition occurs for $p=1$. The model undergoes an Ising and XY transitions for $p=2$ and $p\geq 5$, respectively, and a first-order transition for $p=3$. An XY transition is also expected for $p=4$. Indeed, in the latter case the model obtained for $\kappa=0$ is a generic model with ${\mathbb{Z}}_{4}$ global symmetry, and thus it should have an XY transition, as discussed in Sec. 3.5.2 and, in particular, in footnote 17. The order parameter at these transitions is the gauge-invariant operator $s_{\bm{x}}^{\mathcal{N}}$. This quantity transforms nontrivially only under the ${\mathbb{Z}}_{p}$ subgroup of the global invariance group, so the effective description is given by the LGW Lagrangian (21) with $Q=p$.
From the discussion of the limiting cases it follows that the phase diagram for $p>1$ has the form sketched in Fig. 17, while the phase diagram for $p=1$ is qualitatively similar to the one reported in Fig. 15, with no transition for $\kappa=0$. It is reasonable to assume the transitions along the DD-O line to be analogous to those occurring for $\kappa=0$, while the transitions along the DD-DO are expected to be topological transitions as for $J=0$. Again, the case ${\cal N}=4$ should be discussed separately. Indeed, while the transition for $J=0$ is an Ising gauge transition, for $J\not=0$ we expect an XY gauge phase transition (see the discussion in Sec. 3.5.2 and, in particular, footnote 17).
@@
It is interesting to note that the model with $\mathcal{N}=2$ and any $p\geq 5$ has the same phase diagram and critical behaviors as the $\mathbb{Z}_{2}$-gauge $N$-vector model with $N=2$, which can be obtained by performing a naive limit $q\to\infty$ of the Hamiltonian (84).424242 For $\mathcal{N}=2$ the correlation function $G_{\mathcal{N}}({\bm{x}},{\bm{y}})$ defined in Eq. (85) is equal to $2G_{R}({\bm{x}},{\bm{y}})$ [see Eq. (74)] for $N=2$, if we identify ${\bm{s}}_{\bm{x}}=(\mathrm{Re}\,z_{\bm{x}},\mathrm{Im}\,z_{\bm{x}})$ in Eq. (73). Therefore, for $q\to\infty$ the function $G_{\mathcal{N}}({\bm{x}},{\bm{y}})$ is equivalent to $G_{R}({\bm{x}},{\bm{y}})$ in the $\mathbb{Z}_{2}$-gauge XY model. Thus, for $p\geq 5$ we observe a symmetry enlargement of the global symmetry along the DD-O and DO-O lines. Moreover, for ${\cal N}\geq 4$, correlation functions display an enlarged gauge symmetry along the topological DD-DO transition line.